A smooth sphere A of mass 3m moving on a smooth horizontal table with speed 4u, impiges directly on another smooth sphere B of mass 2m, moving with speed u in the opposite direction to A. The coefficient of restitution between A and B is e.
(a) Find the impulse exerted on A by the impact.
At the moment of the impact, the line of centres of the spheres is perpendicular to a vertical wall which is at a distance x from the point of collision and nearer to B than to A, and B subsequently collides with the wall.
(b) Find in terms of x, the distance of A from the wall at the instant B hits the wall.
1
Expert's answer
2021-02-04T11:43:47-0500

The coefficient of restitution (COR), also denoted by (e), is the ratio of the final to initial relative velocity between two objects after they collide.


"P_1 + P_2 = P_1' + P_2'"

"E_1 + E_2 = E_1' + E_2'"

"3m*4u-2m*u = 3m * x + 2m*y"

"\large\frac{3m*16u^2}{2} + \large\frac{2m*u^2}{2} = \large\frac{3m*x^2}{2}+\large\frac{2m*y^2}{2}"

"{\displaystyle {\text{Coefficient of restitution }}(e)={\frac {\left|{\text{Relative velocity after collision}}\right|}{\left|{\text{Relative velocity before collision}}\right|}}}"

"e = \large\frac{y - x}{4u - u}" "\to y - x = 3ue"

"10mu = m (3x+2y) \to 3x + 2y = 10u"

"48u^2 + 2u^2 = 3x^2 + 2y^2"

"9x^2 + 12xy+4y^2 = 100u^2"

"50u^2 = 3x^2 + 2y^2" / * 2

"9x^2 + 12xy+4y^2 = 6x^2 + 4y^2"

"3x^2 + 12xy = 0"

"3x(x+4y) = 0"

"x = -4y"

"y - x = 5y = 3ue \to y = 0.6ue"

"x = - 4y = -2.4ue" the minus consider the sphere direction

"P_1' = 3mx = 3m * 2.4ue = 7.2mue"